OpenStudy (anonymous):

true or false? 4/5=6/10

jimthompson5910 (jim_thompson5910):

4/5=6/10 4*10=6*5 40 = 30 Since the last equation is false, the first equation is also false.

OpenStudy (anonymous):

We need to get a common denominator. In this case it is 5. Therefore we have\[{4 \over 5} \neq {3 \over 5}\]

OpenStudy (anonymous):

ty